A3 (use Find span of A1, A2, elementary row operations) 1 1 4. = [ 13.43=[1] A =6
Answer #1

considen KIAI + K2 A2 + K3Az = 0 oki + Kat K3 = 0 -Ki tok2 tkz = 0 KI toka tokz okt. kz tkz o. AS ㅕ Katk3 to -k, tkz Kl=0 Kat